OpenStudy (christos):

lim (x ---> -1) (x^2+6x+5)/(x^2-3x-4)

hartnn (hartnn):

can you factor the numerator and denominator ?

OpenStudy (christos):

No :(

hartnn (hartnn):

x^2+6x+5 think of 2 numbers whose product is +5 and sum is +6

OpenStudy (christos):

it gets down to (-1+3)(-1+2)/(-1+1)(-1-4)

OpenStudy (christos):

((-1+3)(-1+2))/((-1+1)(-1-4)) its undetermined

OpenStudy (christos):

@Mertsj

OpenStudy (christos):

@ganeshie8 @ajprincess

OpenStudy (anonymous):

Factorization will work here.. For that you have to first learn how to factorize the things..

OpenStudy (anonymous):

do u need explanation

OpenStudy (raden):

L'Hopital :)

OpenStudy (anonymous):

For numerator part: \[x^2 + 6x + 5\] See here : a = 1 here b= 6 here and c = 5 here by comparing it with \(ax^2 + bx + c\). So, here we have to find factors of 5 that will add to 6 and product to 5. So, by calculations in mind, I am getting +5 and +1. So, this will go like: \[x^2 + 6x + 5 \implies x^2 + x + 5x + 5 \implies x(x + 1) + 5(x + 1) \implies (x + 1)(x + 5)\]

OpenStudy (anonymous):

See, you have found the factors correctly but why aren't you cancelling the common factors?? You got (x+1) on numerator and denominator both, just cancel them and then apply the limits..

OpenStudy (anonymous):

\[\frac{(x + 5)\cancel{(x +1)}}{\cancel{(x + 1)} (x-4)} \implies \frac{(x+5)}{(x-4)}\]

OpenStudy (anonymous):

And now apply the limits here..

OpenStudy (christos):

I see!!

OpenStudy (anonymous):

Oh, you can see?? Great... I can also see...

OpenStudy (christos):

-4/5

OpenStudy (anonymous):

Good..